OpenStudy (hypersniper445):

Please help me I will fan and medal! 1. Graph the equation y = 50x + 200. Give each axis a label. 2. Graph y = 60x

OpenStudy (hypersniper445):

for question 1

OpenStudy (hypersniper445):

for question 2

OpenStudy (hypersniper445):

@triciaal please help

OpenStudy (hypersniper445):

@abb0t @AlexandervonHumboldt2 @arabpride @Ashy98 help me please

OpenStudy (hypersniper445):

does it make sense

OpenStudy (calculusxy):

For the first question, what it's basically asking is that you add a value for x (independent variable or x-coordinate) and then evaluate that function to get a value for y. For instance, the first equation says \(\large y = 50x + 200\). We can start our x value with 0, which will obviously give our y-value of 0 as well (that's just to show the origin of the graph). But more precisely if we have 1 as the x value, what we do is replace x for 1 to get \(\large y = 50(1) + 200\) which is \(\large y = 50 + 200\) or \(\large \color{green}{y = 250}\). As you can see, you have your first coordinate \(\large \color{purple}{(1, 250)}\). Continue doing that same process for every other numbers that you can do. The second one is the same thing. We can have our origin as \(\large (0,0)\). And then if we replace 1 for x, we will have \(\large y= 60(1)\) which is equal to \(\large \color{green}{y = 60}\). So a coordinate for this can be \(\large \color{purple}{(1, 60)}\). Continue doing that same process for every other numbers that you can do.

OpenStudy (hypersniper445):

does the first one look like this

OpenStudy (calculusxy):

Always remember that the x-coordinate will be independent variable and y-axis is dependent variable.

OpenStudy (hypersniper445):

did i do it right

OpenStudy (calculusxy):

Yes

OpenStudy (calculusxy):

But it says to label your axis

OpenStudy (hypersniper445):

oh right

OpenStudy (hypersniper445):

y would be total cost and x would be hours worked

OpenStudy (calculusxy):

Yes that can work. Do you have a specific word problem for this?

OpenStudy (hypersniper445):

yes

OpenStudy (hypersniper445):

Mrs. Solar modeled Contractor 1’s fee formula with the equation y = 50x + 200, where x represents hours worked and y represents total cost. Please use this information for problems #1-7.

OpenStudy (calculusxy):

Yes you're right about the axes labeled.

OpenStudy (hypersniper445):

okay. i was confused on number two because when i graphed it, it seemed to be out of place

OpenStudy (hypersniper445):

do you know what i was doing wrong. the line practically went straight up which didnt make sense

OpenStudy (hypersniper445):

@calculusxy hello?

OpenStudy (calculusxy):

Gimme a moment

OpenStudy (hypersniper445):

okay

OpenStudy (calculusxy):

Can you draw me your second graph so that I can see it?

OpenStudy (hypersniper445):

okay 1 second

OpenStudy (hypersniper445):

i cant recreate it very accurately on a pc but i dont even think this is right anyways

OpenStudy (calculusxy):

You basically start from 0. x-value of 1: 60(1) = 60 x-value of 2, 3, 4 etc. 60(2) = 120 60(3) = 180 60(4) = 240

OpenStudy (calculusxy):

Try to plot those points and see what happens

OpenStudy (hypersniper445):

okay

OpenStudy (hypersniper445):

i dont understand

OpenStudy (calculusxy):

okay wait

OpenStudy (hypersniper445):

do you know what i was doing wrong?

OpenStudy (triciaal):

@HyperSniper445 agree with @calculusxy and gave a medal

OpenStudy (calculusxy):

I don't think that you understood how to evaluate that function with x because that can mislead you to get like the wrong coordinates.

OpenStudy (calculusxy):

But I think you got it now!

OpenStudy (hypersniper445):

okay thank you so much im not very good in math thank you :)

OpenStudy (hypersniper445):

Have a nice day!

OpenStudy (calculusxy):

You're welcome and hope you have a day as well!